The area of a sector of a circle with diameter 12cm is 6.6cm², if the sector is folded to form a cone,
Calculate the radius of the base of the cone.

Answer by mananth(16946)

The area of a sector of a circle with diameter 12cm is 6.6cm², if the sector is folded to form a cone,
Calculate the radius of the base of the cone.
area of a sector of a circle with diameter 12cm is 6.6cm²,
When it is folded to form a cone
The radius of sector becomes the slant height of the cone.
Area of the sector becomes the curved surface area of cone
CSA of cone = pi*r*l
pi*r*6= 6.06
r = 6.06/(6*pi)
r= 0.321 cm